What quantity of a 55% acid solution must be mixed with a 35% solution to produce 600 mL of a 45% solution?
Answer by mananth(16949)

----------------- percent ---------------- quantity
Acid type I 55----------------x ml
Acid type II 35----------------600- x ml
Mixture 45.00% ----------------600

55x+35(600-x)= 45*600

55x+ 21000-35x=27000
55x-35x =27000-21000
20x= 6000
/ 20
x= 300ml 55.00% Acid type I
300ml 35.00% Acid type II
m.ananth@hotmail.ca